# ![[Uncaptioned image]](extracted/6365550/Fig/teaser.png) Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution

### v2

@@ 3.3. Syzygy of Thoughts
CoT | 83.2% | 79.0% | 90.4% | 97.7% | 68.9% | 68.1% | 64.6% | 77.7% | 49.0% |
CoT-SC (n=5) | 86.1% | 81.0% | 93.3% | 98.0% | 71.7% | 70.6% | 66.7% | 80.2% | 52.2% |
SoT (Ours) | 92.1% | 92.5% | 96.1% | 99.2% | 77.2% | 72.3% | 69.1% | 82.5% | 55.0% |
4. Experiments
To evaluate the performance of SoT, we designed a series of experiments, including applicability evaluation (Sec4.1), complex problem reasoning evaluation (Sec4.2), and ablation studies (Sec4.3, 4.4). First, the applicability evaluation tested SoT’s performance across LLMs of different scales, covering mathematical, planning, factual, and logical reasoning, thereby verifying its generalization and stability. Second, the complex problem reasoning evaluation focused on comparing SoT’s superiority in tackling complex tasks on the GSM8K and MATH datasets. Finally, the ablation studies analyzed key components, such as Betti numbers and temperature parameters, to examine their impact on the complexity, expressiveness, and robustness of the reasoning chain. These experiments comprehensively validate the theoretical and practical value of SoT.
4.1. Applicability Evaluation
The reasoning performance of chain-of-thought methods (such as CoT and CoT-SC) varies significantly across models of different sizes, especially on lightweight models. We aim to validate whether SoT can consistently enhance the reasoning capabilities of various LLMs and demonstrate its generalization and stability across models and task scenarios. For this purpose, we applied CoT, CoT-SC (n=5), and SoT to each model, with the default Betti number set to 7 and mapping set to 3, recording the accuracy across nine datasets.
Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3. We selected five different scales of LLMs, including GPT-4o-mini (lightweight), Qwen2.5-Coder-7B-Instruct (7B), Gemma-3-12b-it (12B), Gemma-3-27b-it (27B), and Qwen2.5-VL-72B-Instruct (72B), covering models from lightweight to large-scale. The baseline models were standard CoT and CoT-SC (n=5). The test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH (multi-task answering), Data (date prediction), and CLUTRR (logical reasoning), covering nine datasets in total.
Experimental Results: The results are shown in Table 1. SoT significantly outperforms the baselines across all models and datasets. For example, on GSM8K with GPT-4o-mini, SoT achieved 96.0%, improving by 10.9% over CoT (85.1%) and 5.9% over CoT-SC (90.1%). On the logical reasoning task CLUTRR, SoT achieved 75.7%, surpassing CoT (65.9%) and CoT-SC (72.4%). On Qwen2.5-VL-72B-Instruct, SoT achieved 89.4% on AQUA, outperforming CoT (81.1%) and CoT-SC (83.9%). For the lightweight model Qwen2.5-Coder-7B-Instruct, SoT improved GSM8K accuracy to 89.1%, 11.9% higher than CoT (77.2%), showing its superiority in resource-constrained scenarios. Furthermore, on Gemma-3-12b-it, SoT achieved 92.5% on SVAMP, close to the 95.8% achieved by Gemma-3-27b-it, indicating that its generalization ability is not limited by model scale. This result demonstrates the significant advantage of SoT’s structured design in improving reasoning performance and stability across models.
4.2. Complex Reasoning Evaluation
@@
To compare the complex problem reasoning capabilities of SoT with mainstream methods, we evaluated SoT and compared it with extensive open-source data from various papers (Chen et al., 2025) . Specifically, we applied the SoT reasoning chain to GPT-4o-mini and tested its reasoning accuracy on the GSM8K and MATH datasets, while comparing it with other methods from the literature.
@@
Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3. The baselines included CoT and various popular chain-of-thought variants (such as MathPrompter, QuaSAR, MathDivide, etc.). The datasets used were GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
@@
No-CoT (Deng et al., 2024) | Mistral-7B | 38.0% | - |
ICoT-SI (Deng et al., 2024) | Mistral-7B | 51.0% | - |
- | RecurrentBlock-3.5B | 42.1% | - |
MathCoder-CL (Wang et al., 2023b) | Code-Llama-7B | 67.8% | 30.2% |
MAmmoTH (Yue et al., 2023) | Code-Llama-7B | 59.4% | - |
Brain (Chen et al., 2024a) | Code-Llama-7B | 74.0% | - |
SQ-VAE (Wang et al., 2023a) | Llama-2-7B | 40.0% | 7.0% |
Self-Rewarding (Chen et al., 2024b) | Llama-2-7B | 40.0% | 10.7% |
STaR (Zelikman et al., 2022) | Llama-2-7B | 58.2% | 16.0% |
ENVISIONS (Xu et al., 2024) | Llama-2-7B | 59.0% | 19.0% |
MetaMath (Yu et al., 2023b) | Llama-2-7B | 66.5% | - |
ToRA-Code (Gou et al., 2023) | Llama-2-7B | 72.6% | - |
OVM (Yu et al., 2023a) | Llama-2-7B | 73.7% | - |
- | Llama-3.1-8B | 56.7% | 20.3% |
- | Llama-3.1-70B | 85.5% | 41.4% |
- | Llama-3.1-405B | 89.0% | 53.8% |
- | NuminaMath-7B-CoT | 75.4% | 55.2% |
- | DeepSeek-Coder-7B | 77.4% | 44.4% |
- | Qwen2-7B | 79.9% | 44.2% |
- | Qwen2-Math-7B | 80.4% | 50.4% |
SIaM (Yu et al., 2024) | Qwen-2-Math-Base | 81.5% | 50.0% |
- | Internlm2-math-plus-7B | 84.0% | 54.4% |
OMI2 (Li et al., 2025) | Qwen2.5-Coder-7B | 84.1% | 72.3% |
CODEI/O++ (Li et al., 2025) | Qwen2.5-Coder-7B | 85.7% | 72.1% |
PyEdu (Li et al., 2025) | Qwen2.5-Coder-7B | 85.8% | 71.4% |
CODEI/O (Li et al., 2025) | Qwen2.5-Coder-7B | 86.4% | 71.9% |
OC-SFT-1 (Li et al., 2025) | Qwen2.5-Coder-7B | 86.7% | 70.9% |
WI (Li et al., 2025) | Qwen2.5-Coder-7B | 87.0% | 71.4% |
WI (Full) (Li et al., 2025) | Qwen2.5-Coder-7B | 87.0% | 71.1% |
OMI2 (Full) (Li et al., 2025) | Qwen2.5-Coder-7B | 88.5% | 73.2% |
- | DeepSeekMath-7B-RL | 88.2% | 51.7% |
CoMAT (Leang et al., 2024) | GPT-4 | 93.7% | - |
CoT (Ranaldi et al., 2025) | GPT-4 | 94.5% | - |
FCoT (Lyu et al., 2023) | GPT-4 | 95.0% | - |
MathPrompter (Imani et al., 2023) | GPT-4 | 95.6% | - |
QuaSAR (Radford et al., 2018) | GPT-4 | 96.5% | - |
MathDivide (Srivastava and Gandhi, 2024) | GPT-4 | 96.8% | - |
@@
Experimental Results: As shown in Table 2, SoT achieved 96.0% on GSM8K and 79.1% on MATH on GPT-4o-mini. On GSM8K, SoT’s performance is close to the best result achieved by GPT-4 (e.g., MathDivide’s 96.8%) (Chen et al., 2025) and outperforms the best 7B model (e.g., OMI2 Full’s 88.5%) (Chen et al., 2025). On MATH, SoT’s 79.1% significantly outperforms the best mainstream 7B models (e.g., OMI2 Full’s 73.2%). This result indicates that SoT can achieve reasoning capabilities close to those of closed-source large models on lightweight models, significantly narrowing the performance gap between open-source models and GPT-4, validating its superiority in complex mathematical reasoning.
4.3. Betti Number Sensitivity
The SoT reasoning chain regulates topological constraints through the Betti number, and its value directly affects performance. To investigate the optimal configuration of the Betti number and its dynamic impact on structural expressiveness and regularization effects, we designed ablation experiments to reveal its key role in optimizing the reasoning process. For this, we systematically adjusted the Betti number and recorded accuracy changes under different mapping configurations. The goal of the experiment was to analyze the impact of the Betti number on the reasoning chain’s expressiveness and identify the performance saturation point. The results are presented in Figure 5, showing the trend of accuracy changes with varying Betti numbers, with particular attention to its non-monotonic properties and optimal configuration.
Experimental Configuration: The experiment was based on GPT-4o-mini and conducted on the GSM8K dataset. The Betti number was incremented from 1 to 10, with mapping set to 1, 2, and 3 to evaluate performance under different topological decomposition granularities.
Experimental Results: As shown in Figure 5, accuracy changed non-monotonically as the Betti number increased under mapping configurations of 1, 2, and 3. Specifically, when the Betti number started from 1, accuracy improved significantly, indicating that a small number of topological constraints effectively enhanced structural expressiveness. However, as the Betti number increased further, the performance gain gradually diminished and stabilized. The optimal saturation point was reached at 7, beyond which there was no noticeable improvement. The results reveal the dynamic regulatory effect of the Betti number in reasoning chain modeling, and that moderate topological constraints are crucial for optimizing SoT’s structured design, providing strong support for its theoretical foundation.
4.4. Stability Analysis
The temperature parameter influences the diversity of content generated by LLMs, which may challenge reasoning stability. We aimed to explore whether the SoT method could maintain its performance advantage under varying temperature conditions, thereby validating the adaptability of its structured design. To this end, we systematically evaluated SoT and CoT under different temperature conditions, focusing on the stability of accuracy and the relationship between reasoning diversity and consistency. The experiment involved adjusting the temperature step by step and observing the performance trends of the two methods under changing diversity.
Experimental Configuration: The experiment was based on GPT-4o-mini, with the Betti number fixed at 7 and mapping set to 3. The nine test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH, Date (multi-task QA), and CLUTRR (logical reasoning). The temperature parameter was adjusted from 0.0 to 1.0 in steps of 0.1, and the performance of SoT and CoT was compared. Five reasoning paths were generated in each experiment, and the average accuracy was recorded to reduce the impact of randomness.
Experimental Results:As shown in Figure 6, the accuracy of SoT across multiple datasets varies minimally with temperature (from 0.0 to 1.0), exhibiting stability with small amplitude. Figure 8indicates that the accuracy of CoT fluctuates significantly within the same temperature range, with the amplitude increasing as the temperature rises, especially at high temperatures (close to 1.0), where the fluctuations become more pronounced. The boxplot in Figure 7further compares the accuracy distributions of both methods across datasets: the box of SoT is relatively small, with few outliers, and its accuracy is concentrated with high consistency across datasets, even at high temperatures, without significant outliers. In contrast, the box of CoT is larger, with a more scattered distribution and more outliers, especially at high temperatures, where the reasoning consistency diminishes. Analysis suggests that the high temperature-induced diversity in CoT weakens logical coherence, while SoT maintains reasoning stability.
5. conclusion
In this paper, we introduce the SoT reasoning framework, integrating the MFR principle from algebraic geometry and homological algebra into the CoT paradigm, significantly enhancing LLM performance in complex reasoning tasks. SoT addresses traditional CoT limitations in high-dimensional, nonlinear, and logical scenarios by employing a structured decomposition strategy—utilizing Module, Freeness, Mapping, Exactness, Minimality, and Betti numbers—to create compact, logically consistent reasoning units.
While SoT has shown excellent reasoning capabilities with LLMs, further exploration is needed. We plan to extend the topological decomposition to multimodal reasoning, including inputs like images and tables, to verify its adaptability in cross-modal tasks. Additionally, incorporating iterative MFR concepts will enhance step-by-step problem refinement, optimizing reasoning paths dynamically and improving SoT’s efficiency and accuracy in multimodal and high-dimensional tasks.

### v3

@@ 3.2 Syzygy of Thoughts
| CoT (Wei et al., 2022) | 83.5% | 79.3% | 82.7% | 91.0% | 69.2% | 68.4% | 64.9% | 78.0% | 49.3% |
| CoT-SC (Wang et al., 2022b) | 85.9% | 80.8% | 85.1% | 93.2% | 71.5% | 70.4% | 66.5% | 80.0% | 52.0% |
| GoT (Besta et al., 2024) | 89.2% | 86.8% | 86.8% | 93.9% | 74.5% | 71.6% | 68.0% | 81.5% | 53.7% |
| AoT (Teng et al., 2025) | 91.4% | 91.8% | 87.5% | 93.5% | 76.7% | 71.9% | 68.7% | 82.0% | 54.8% |
| SoT (Ours) | 92.5% | 92.9% | 88.5% | 94.6% | 77.6% | 72.7% | 69.5% | 82.9% | 55.4% |
4 Experiments
4.1 Applicability Evaluation
Configuration. As shown in Table 1, we compare the proposed SoT method with four representative reasoning strategies: Chain-of-Thought (CoT) (Wei et al., 2022), Self-Consistency CoT (CoT-SC) (Wang et al., 2022b), Graph-of-Thought (GoT) (Besta et al., 2024), and Atom-of-Thought (AoT) (Teng et al., 2025). The evaluation covers nine benchmarks spanning four categories: mathematical word problems (GSM8K, SVAMP, MultiArith, ASDiv, AQUA), general-knowledge and multitask QA (MMLU, BBH), temporal reasoning (Date), and relational logical reasoning (CLUTRR). For CoT-SC, we follow the standard setting with sampled reasoning paths. GoT is instantiated with beam size and search depth , while AoT uses a maximum search depth of , consistent with the original works. We instantiate all methods on five backbone LLMs: GPT-4o-mini, Qwen2.5-Coder-7B-Instruct, Qwen2.5-VL-72B-Instruct, Gemma-3-27b-it, and Gemma-3-12b-it. For each backbone, we keep decoding hyper-parameters (e.g., temperature, top-, maximum generation length) fixed across methods and only change the reasoning protocol. We report mean accuracy and standard deviation over multiple random seeds to account for sampling variance.
Results. From Table 1, SoT consistently achieves the best performance across all benchmarks and all backbone models. On the mathematical reasoning benchmarks, SoT improves over the strongest baseline AoT by around percentage point on GSM8K and SVAMP for every model family (e.g., on GPT-4o-mini, GSM8K accuracy increases from to ), while also yielding smaller but steady gains on MultiArith and ASDiv. On the more challenging AQUA dataset, SoT brings additional improvements of roughly – points across different backbones, indicating that searching over structured thoughts helps the model better handle multi-step algebraic reasoning and eliminate distractor options. For general-knowledge and multitask reasoning, SoT further improves MMLU and BBH by approximately – points over AoT on all models, with Qwen2.5-VL-72B and Gemma-3-27b-it achieving the highest absolute accuracies. Finally, on temporal and relational logical reasoning (Date and CLUTRR), SoT again attains the best results under every backbone, with gains of up to about point, suggesting that the structured search procedure is beneficial beyond pure arithmetic reasoning. Overall, SoT dominates all baselines in every setting in Table 1 while maintaining comparable or even lower variance, demonstrating that it provides a stable and broadly applicable improvement to LLM reasoning.
4.2 Cost Test
To systematically compare the trade-offs between accuracy and search cost across different reasoning frameworks, we construct the three-dimensional bubble plot shown in Figure 4, which places five methods under a unified lens: CoT, CoT-SC, GoT, AoT, and our proposed SoT. All methods are evaluated using the same model and the same set of nine benchmark tasks. In the plot, the vertical axis represents the macro-average accuracy over the nine tasks, the horizontal axis denotes the average number of candidate reasoning paths generated per problem, and the bubble area is proportional to the average number of tokens consumed per problem. We provide the detailed definition of our path-counting metric in Appendix 4.2.
Based on the above definitions, we visualize a representative set of method configurations. Along the horizontal axis, the number of candidate paths increases from left to right; along the vertical axis, the macro-average accuracy improves from bottom to top; and the bubble size grows with the token cost. When the number of candidate paths and the token budget are comparable, SoT attains substantially higher macro-average accuracy than CoT-SC, GoT, and AoT. Moreover, as the number of syzygies increases from to , the marginal gains diminish noticeably—indicating that SoT achieves near-saturated performance even with only a small number of syzygies. Overall, SoT occupies a particularly favorable region in the three-way trade-off space: with a moderate number of candidate paths and an acceptable token cost, it achieves the highest macro-average accuracy and significantly enhances the effectiveness of each individual candidate path through its syzygy-based structure.
4.3 Abalation Study
Betti Number Sensitivity. The SoT reasoning chain regulates topological constraints through the Betti number, and its value directly affects performance. To investigate the optimal configuration of the Betti number and its dynamic impact on structural expressiveness and regularization effects, we designed ablation experiments to reveal its key role in optimizing the reasoning process. For this, we systematically adjusted the Betti number and recorded accuracy changes under different mapping configurations. The goal of the experiment was to analyze the impact of the Betti number on the reasoning chain’s expressiveness and identify the performance saturation point. The results are presented in Figure 5, showing the trend of accuracy changes with varying Betti numbers, with particular attention to its non-monotonic properties and optimal configuration.
Configuration. The experiment was based on GPT-4o-mini and conducted on the GSM8K dataset. The Betti number was incremented from 1 to 10, with mapping set to 1, 2, and 3 to evaluate performance under different topological decomposition granularities.
Results. As shown in Figure 5 (left), accuracy changed non-monotonically as the Betti number increased under mapping configurations of 1, 2, and 3. Specifically, when the Betti number started from 1, accuracy improved significantly, indicating that a small number of topological constraints effectively enhanced structural expressiveness. However, as the Betti number increased further, the performance gain gradually diminished and stabilized. The optimal saturation point was reached at 7, beyond which there was no noticeable improvement. The results reveal the dynamic regulatory effect of the Betti number in reasoning chain modeling, and that moderate topological constraints are crucial for optimizing SoT’s structured design, providing strong support for its theoretical foundation.
Stability Analysis. The temperature parameter influences the diversity of content generated by LLMs, which may challenge reasoning stability. We aimed to explore whether the SoT method could maintain its performance advantage under varying temperature conditions, thereby validating the adaptability of its structured design. To this end, we systematically evaluated SoT and CoT under different temperature conditions, focusing on the stability of accuracy and the relationship between reasoning diversity and consistency. The experiment involved adjusting the temperature step by step and observing the performance trends of the two methods under changing diversity.
Configuration. The experiment was based on GPT-4o-mini, with the Betti number fixed at 7 and mapping set to 3. The nine test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH, Date (multi-task QA), and CLUTRR (logical reasoning). The temperature parameter was adjusted from 0.0 to 1.0 in steps of 0.1, and the performance of SoT and CoT was compared. Five reasoning paths were generated in each experiment, and the average accuracy was recorded to reduce the impact of randomness.
Results. As shown in Figure 5 (right), SoT’s accuracy remains stable across temperature variations (0.0 to 1.0) with minimal fluctuation. In contrast, CoT’s accuracy fluctuates significantly, especially at higher temperatures, where the amplitude increases. The boxplot in Figure 5 (middle) reveals that SoT’s accuracy is concentrated with little variation and few outliers across datasets, even at high temperatures. On the other hand, CoT shows a wider spread with more outliers, especially at higher temperatures, where reasoning consistency decreases. This analysis indicates that CoT’s high temperature-induced diversity weakens logical coherence, while SoT maintains stability.
5 Conclusions
We introduced SoT, an MFR-inspired test-time reasoning framework that reframes CoT from a single linear trace into a syzygy-structured family of interdependent reasoning paths. By mapping MFR notions to concrete prompting and selection operations, SoT explicitly organizes auxiliary conditions, constructs candidate “syzygies,” verifies logical closure, and prunes redundancy to produce compact yet faithful solutions. Across nine benchmarks spanning math, general knowledge, multitask QA, temporal reasoning, and relational logic, and across five LLM backbones, SoT consistently outperforms strong baselines while maintaining comparable or lower variance, indicating both improved accuracy and more stable inference. Moreover, SoT occupies a favorable region in the accuracy–cost–paths trade-off, achieving near-saturated gains with only a small number of syzygies, and exhibiting strong robustness under sampling diversity.
6 Limitations
While SoT demonstrates strong reasoning capabilities, further research is needed. We plan to extend the topological decomposition to multimodal reasoning, including images and tables, to test its adaptability in cross-modal tasks. Additionally, incorporating iterative MFR concepts will refine problem-solving steps, optimize reasoning paths, and improve SoT’s efficiency in multimodal and high-dimensional tasks.
Appendix A Appendix
A.1 Full Pseudocode of the SoT Pipeline
In this appendix, we present the full pseudocode of the SoT pipeline. Algorithm 1 provides an algorithmic implementation of the LLM-based minimal free resolution discussed in Section 3, turning the conceptual modules in Figure 3 and the abstract diagram in Figure 6 into an executable procedure.
Explanation of Algorithm 1.
We briefly explain the main components of Algorithm 1, which instantiates the SoT framework described in Section 3. Analyse and homological complexity. Given a reasoning problem and its metadata , the subroutine Analyse estimates the regularity and the projective dimension . These quantities summarize the Betti-number profile introduced in Section 3: determines the number of auxiliary (freeness) conditions, and specifies how many candidate resolutions (syzygies) are explored. Larger values of indicate higher problem complexity. Generating auxiliary conditions (Freeness). The loop over constructs a set of auxiliary conditions, corresponding to the “Freeness” module in Figure 3. Each call to LLMGenerateAuxCondition conditions on previously created free modules and maps to propose a new constraint . These constraints guide the LLM toward structurally coherent reasoning decompositions. Constructing candidate syzygies (Syzygy of Thoughts). The loop over implements the “Syzygy of Thoughts” mechanism from Section 3. Each call to LLMResolve yields a candidate reasoning chain that satisfies the auxiliary conditions. From an algebraic viewpoint, each functions as a syzygy relating different generators of , and the list collects all such candidates. Selecting a minimal resolution (Minimality). Once the candidate set is formed, the subroutine ArgMinScore selects the best resolution according to a scoring function . This function, implemented via an LLM scoring prompt, evaluates both structural regularity (e.g., consistency with freeness and mapping constraints) and task-specific correctness. This step aligns with the “Minimality” module in Figure 3. Extracting the final answer (Exactness). Finally, ExtractFinalAnswer converts the selected reasoning chain into the final task-specific answer. Pattern matching techniques (e.g., regular expressions or JSON parsing) ensure the output is consistent with the resolved structure and adheres to benchmark formatting requirements, realizing the “Exactness” principle.
A.2 Conceptual Illustration of the SoT Pipeline
Explanation.
Figure 6 provides a concrete walk-through of the SoT framework using a simple arithmetic problem. The illustration makes explicit how SoT transforms a natural-language question into a structured multi-stage reasoning process.
(1) Module and Betti Numbers. The LLM first rewrites the question into a normalized “module” form and performs a lightweight complexity analysis. Based on the estimated problem structure, it predicts the required number of auxiliary conditions, corresponding to the Betti-number estimate discussed in Section 3. (2) Freeness. The LLL generates non-redundant auxiliary conditions (e.g., price per pair, number of children, number of pairs per child). These auxiliary facts form the initial free modules , which constrain the search space of valid reasoning chains. (3) Mappings. The LLM organizes auxiliary conditions into candidate reasoning strategies by combining freeness conditions in different orders. Each such combination corresponds to a candidate syzygy, i.e., a structured reasoning chain that reflects different possible mappings between the modules in the minimal free resolution. (4) Exactness. The LLM inspects each candidate strategy and verifies whether every reasoning step is justified, gap-free, and based on explicitly stated assumptions. This matches the exactness requirement in the algebraic SoT formulation. (5) Minimality. Among all valid strategies, the LLM selects the one with the smallest reasoning footprint, balancing accuracy, complexity, and simplicity. This corresponds to the minimality principle described in Section 3. (6) Final Answer. Once the optimal syzygy is chosen, the LLM extracts the final numerical answer while ensuring consistency with all earlier reasoning stages. This mirrors the final extraction step in Algorithm 1.
Relation to Algorithm 1.
This example illustrates the same pipeline implemented formally in Algorithm 1: Analyse corresponds to Step 1; LLMGenerateAuxCondition and LLMResolve correspond to Steps 2–3; ArgMinScore corresponds to Step 4; and ExtractFinalAnswer corresponds to Step 5. The illustration provides an intuitive view complementary to the algorithmic description.
A.3 Complex Reasoning Evaluation
@@
To compare the complex problem reasoning capabilities of SoT with mainstream methods, we evaluated SoT and compared it with extensive open-source data from various papers Chen et al. (2025). Specifically, we applied the SoT reasoning chain to GPT-4o-mini and tested its reasoning accuracy on the GSM8K and MATH datasets, while comparing it with other methods from the literature.
@@
Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3. The baselines included CoT and various popular chain-of-thought variants (such as MathPrompter, QuaSAR, MathDivide, etc.). The datasets used were GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021).
@@
| No-CoT Deng et al. (2024) | Mistral-7B | 38.0% | - |
| ICoT-SI Deng et al. (2024) | Mistral-7B | 51.0% | - |
| - | RecurrentBlock-3.5B | 42.1% | - |
| MathCoder-CL Wang et al. (2023a) | Code-Llama-7B | 67.8% | 30.2% |
| MAmmoTH Yue et al. (2023) | Code-Llama-7B | 59.4% | - |
| Brain Chen et al. (2024a) | Code-Llama-7B | 74.0% | - |
| SQ-VAE Wang et al. (2023b) | Llama-2-7B | 40.0% | 7.0% |
| Self-Rewarding Chen et al. (2024b) | Llama-2-7B | 40.0% | 10.7% |
| STaR Zelikman et al. (2022) | Llama-2-7B | 58.2% | 16.0% |
| ENVISIONS Xu et al. (2024) | Llama-2-7B | 59.0% | 19.0% |
| MetaMath Yu et al. (2023b) | Llama-2-7B | 66.5% | - |
| ToRA-Code Gou et al. (2023) | Llama-2-7B | 72.6% | - |
| OVM Yu et al. (2023a) | Llama-2-7B | 73.7% | - |
| - | Llama-3.1-8B | 56.7% | 20.3% |
| - | Llama-3.1-70B | 85.5% | 41.4% |
| - | Llama-3.1-405B | 89.0% | 53.8% |
| - | NuminaMath-7B-CoT | 75.4% | 55.2% |
| - | DeepSeek-Coder-7B | 77.4% | 44.4% |
| - | Qwen2-7B | 79.9% | 44.2% |
| - | Qwen2-Math-7B | 80.4% | 50.4% |
| SIaM Yu et al. (2024) | Qwen-2-Math-Base | 81.5% | 50.0% |
| - | Internlm2-math-plus-7B | 84.0% | 54.4% |
| OMI2 Li et al. (2025) | Qwen2.5-Coder-7B | 84.1% | 72.3% |
| CODEI/O++ Li et al. (2025) | Qwen2.5-Coder-7B | 85.7% | 72.1% |
| PyEdu Li et al. (2025) | Qwen2.5-Coder-7B | 85.8% | 71.4% |
| CODEI/O Li et al. (2025) | Qwen2.5-Coder-7B | 86.4% | 71.9% |
| OC-SFT-1 Li et al. (2025) | Qwen2.5-Coder-7B | 86.7% | 70.9% |
| WI Li et al. (2025) | Qwen2.5-Coder-7B | 87.0% | 71.4% |
| WI (Full) Li et al. (2025) | Qwen2.5-Coder-7B | 87.0% | 71.1% |
| OMI2 (Full) Li et al. (2025) | Qwen2.5-Coder-7B | 88.5% | 73.2% |
| - | DeepSeekMath-7B-RL | 88.2% | 51.7% |
| CoMAT Leang et al. (2024) | GPT-4 | 93.7% | - |
| CoT Ranaldi et al. (2025) | GPT-4 | 94.5% | - |
| FCoT Lyu et al. (2023) | GPT-4 | 95.0% | - |
| MathPrompter Imani et al. (2023) | GPT-4 | 95.6% | - |
| QuaSAR Radford et al. (2018) | GPT-4 | 96.5% | - |
| MathDivide Srivastava and Gandhi (2024) | GPT-4 | 96.8% | - |
@@
Experimental Results: As shown in Table 2, SoT achieved 96.0% on GSM8K and 79.1% on MATH on GPT-4o-mini. On GSM8K, SoT’s performance is close to the best result achieved by GPT-4 (e.g., MathDivide’s 96.8%) Chen et al. (2025) and outperforms the best 7B model (e.g., OMI2 Full’s 88.5%) Chen et al. (2025). On MATH, SoT’s 79.1% significantly outperforms the best mainstream 7B models (e.g., OMI2 Full’s 73.2%). This result indicates that SoT can achieve reasoning capabilities close to those of closed-source large models on lightweight models, significantly narrowing the performance gap between open-source models and GPT-4, validating its superiority in complex mathematical reasoning.
A.4 Path Counting and Token-Cost Metrics
To align the notion of “paths” across different methods, we adopt a unified definition. For a method and a sample , let denote the number of candidate reasoning trajectories generated on that contain a complete chain of reasoning and yield a directly scorable answer. The horizontal axis in our plots reports
which reflects how many full candidate solutions, on average, the method explores per problem.
Under this formulation:
CoT: Each problem produces exactly one chain-of-thought, hence .
CoT-SC(): Each problem samples independent CoT trajectories followed by majority voting, thus .
GoT(): During multi-round graph search, each invocation that produces a complete candidate solution is counted as a path. Internal graph-expansion calls that generate intermediate nodes without a final answer do not count as paths, though their token usage is included in the bubble size.
AoT(): Along the Markov chain, solving each of produces a directly scorable candidate answer. Hence each solve() call is counted as one path, and approximates the average number of solver calls per Markov chain.
SoT(): In SoT, LLMRESOLVE generates syzygies , each constituting a full structured reasoning chain capable of producing an answer. Therefore the number of candidate paths is exactly the number of syzygies:
The auxiliary calls used during the freeness stage do not count as paths, but all of their token usage is included in the overall cost.
The bubble area in the figure corresponds to the average tokens per problem, computed as the total number of tokens consumed by all LLM calls (generation, intermediate analysis, candidate solutions, and scoring), averaged over the evaluation set. This metric reflects the overall reasoning cost of each method.